OpenStudy (kaynick):

Graph the line. Y=-1/5x-4

OpenStudy (greatlife44):

\[y = -\frac{ 1 }{ 5 }x-4\]

OpenStudy (greatlife44):

to graph this line here's what you do. you can insert an x value and you'll get a y value. let's plug in x = 0 \[y = -\frac{ 1 }{ 5 }(0)-4\] \[y = -4, when~x = 0\]